OpenStudy (anonymous):

4and three thenths minus 3 and 13 over 15 ?

OpenStudy (saifoo.khan):

wht 'and' ?

OpenStudy (anonymous):

lol, you can't read saifoo?

OpenStudy (saifoo.khan):

i can, thats y i wrote! :P

OpenStudy (anonymous):

4 3/ 10 - 3 13/15

OpenStudy (saifoo.khan):

13/30

OpenStudy (saifoo.khan):

hero, chk out christie's ques!

OpenStudy (anonymous):

how did you do it I got 15/58

OpenStudy (anonymous):

I did already saifoo

OpenStudy (saifoo.khan):

idk, my calc got it!

OpenStudy (saifoo.khan):

lol.

OpenStudy (anonymous):

ok what abt 6 7/10 - 4 9/10 ?

OpenStudy (saifoo.khan):

9/5

OpenStudy (anonymous):

its improper?

OpenStudy (saifoo.khan):

yes

OpenStudy (saifoo.khan):

or u can write: 1 4/5

OpenStudy (anonymous):

ok 16 1/8 - 8 3/8 (;

OpenStudy (anonymous):

?

OpenStudy (anonymous):

4and three thenths minus 3 and 13 over 15\[\left(4+\frac{3}{10}\right)-\left(3+\frac{13}{15}\right) =\frac{43}{10}-\frac{58}{15}=\frac{13}{30} \]

OpenStudy (anonymous):

ok 16 1/8 - 8 3/8

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[\left(16+\frac{1}{8}\right)-\left(8+\frac{3}{8}\right)=\frac{31}{4} \]

OpenStudy (anonymous):

ok 9 3/7 - 4 6/7

OpenStudy (anonymous):

\[\frac{66}{7}-\frac{34}{7}=\frac{32}{7} \]

OpenStudy (anonymous):

21 9/15 - 6 14/15

OpenStudy (anonymous):

\[\frac{248}{15} \]

OpenStudy (anonymous):

thats improper tho what is it as a fraction?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[16+\frac{8}{15} \]

OpenStudy (anonymous):

I got 14 2/3

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

???//

OpenStudy (anonymous):

Are you asking is 14 2/3 correct for 248/15 ?

OpenStudy (anonymous):

yes (:

OpenStudy (anonymous):

To convert 248 to a fraction first divide 248 by 15. What quotient do you get?

OpenStudy (anonymous):

for the answer I got 14 2/3

OpenStudy (anonymous):

248 divided by 15 equals 16 plus a remainder. Can you divide 248 by 15 either by hand or using a calculator?

OpenStudy (anonymous):

ok i got 15 2/3

OpenStudy (anonymous):

\[15+2/3=\frac{47}{3} \] not 248/15